OpenStudy (mollykenney):

PLEASE HELP!! Use the Change of Base Formula to evaluate each expression. log26 5

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is the base 2 or 26?

OpenStudy (misty1212):

\[\log_2(65)\] or \[\log_{26}(5)\]

OpenStudy (misty1212):

in any case this is what you need \[\log_b(A)=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (mollykenney):

the 2nd

OpenStudy (misty1212):

ok so \[\frac{\ln(5)}{\ln(26)}\] and a calculator

OpenStudy (misty1212):

since you are on a computer, you do not need a calculator you can use this http://www.wolframalpha.com/input/?i=ln%285%29%2Fln%2826%29

OpenStudy (mollykenney):

ah thank you so much misty

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]